OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -1 + 2 + 5 + 8 + ... + 44

OpenStudy (john_es):

You see that the distance between each term is, 3, \[5-2=3\]So, you can do the following sum, \[\sum_{j=1}^{15}(-1+3j)\] That is the answer. You know that you must sum until j=15 because, \[44=-1+3j\Rightarrow j=15\]

OpenStudy (anonymous):

is this the same as your answer: